OpenStudy (anonymous):

The lengths of the legs of a right triangle are 25 cm and 60 cm. What is the length of the hypotenuse? A. 9.2 centimeters B. 65 centimeters C. 35 centimeters D. 42.5 centimeters

OpenStudy (♪chibiterasu):

Use the Pythagorean Theorom: \[a^2 + b^2 = c^2\]

OpenStudy (cwrw238):

Hypotenuse squared = 25^2 + 60^2 work that out and take the aquare root of the answer this is pythagoras theorem

OpenStudy (♪chibiterasu):

Yeah, that. Replace the gien values with a and b (the order doesn't matter) and solve.

OpenStudy (♪chibiterasu):

given*

OpenStudy (anonymous):

Ok thanks everyone c:

OpenStudy (cwrw238):

yw